# Multi-Source Urban Traffic Flow Forecasting with Drone and Loop Detector Data

### v1

@@ III-C Dataset Creation
Although existing drone datasets have already provided high-quality vehicle trajectory data, e.g., pNEUMA [10] and highD [44], the aim of this work is to predict the traffic states for all road segments within a large urban region, which is beyond the scope of available datasets. For this reason, detailed vehicle trajectory data are generated using the microscopic traffic simulator Aimsun and its road network for central Barcelona [45]. The simulation environment has been properly calibrated based on real-world traffic scenarios, and it has been applied in various analyses on urban traffic, such as network wide-perimeter control [46, 47]. As shown in Figure 3, the network consists of many road segments, intersections as well as centroids ($\odot$ like) that serve as origins and destinations of vehicles. The fine details from microscopic traffic simulations allow great flexibility for subsequent processing and analysis, e.g., the computation of vehicle flow, density and speed. Concretely, there are mainly three types of available information:
  • 1.
@@
When computing traffic variables from trajectories, the characteristics of different sensors need to be carefully considered to resemble the real-world data collection process. In this work, we focus on the speed measurements from two types of sensors: inductive loop detectors and drones. Inductive loop detectors can only measure vehicle speeds at certain points, as they are installed at fixed locations. Since vehicles can change their speeds within a road segment, the speed measurements of a loop detector are valid only in its proximity and often cannot represent the situation of the whole segment. On the contrary, drones can capture nearly all vehicles in their field of view, therefore it is possible to compute accurate segment-level speeds by using the complete vehicle trajectories. In this paper, we refer to these two types of speeds as point speed and segment speed respectively, and both of them can be extracted with the trajectory data from microscopic simulation.
@@
For a loop detector installed $d$ meters from the start, if a trajectory split has $x_{s}<d<x_{e}$, then the corresponding vehicle is detected at the timestep of the trajectory split. We count the number of detected vehicles $n$ in a time interval (e.g., every 3 minutes), and arithmetically average their speeds to get point speed as
@@
By choosing different time resolutions for Equation 6 and 7, the simulated loop detector and drone measurements can be computed individually for each road segment. Similarly, regional speed can be obtained by applying Equation 6 to all the trajectory splits in a region during a time interval.
Then, the training samples for the prediction model can be generated using a sliding window approach as depicted in Figure 5. Each simulation session lasts for 4 hours in the simulation world, with a 15-minute warm-up and 1 hour 45 min main demand, while the real-world run time differs according to the traffic demand. This demand profile is intended to simulate the morning peak hours in the road network. The sliding window will start at the end of the warm-up period up to 2 hours after the beginning of the main demand, and the window size is 1 hour with a 3-minute step size between consecutive samples. Inside a window, the first half is used as input and the second half is used as targeted output for the prediction model. Therefore each simulation session will have 20 training samples, each with 30-minute input and 30-minute output.
@@ IV Experiments
This section will present the dataset description at both micro and macro levels, implementation details, and the evaluation of prediction performance.
Generally, SimBarca introduces two input modalities, i.e., vehicle speeds from loop detectors and drone monitoring, and two prediction tasks, i.e., road segment speed prediction and regional speed prediction. As depicted in Figure 5, the prediction model is required to forecast the future 30 minutes given the data of the past 30 minutes. The input modalities provide measurements of road segments with different time resolutions, where loop detectors report speeds every 3 minutes and drones report speeds every 5 seconds. Both prediction tasks are labeled with future speed values every 3 minutes. The segment speeds computed by Equation 7 are used as segment-level labels, and the regional labels are obtained by aggregating all trajectory splits within a region and applying the same formula. When computing the point speed, the loop detector is assumed to be installed at the midpoint of a road segment.
Although in reality, drones are able to provide high-resolution video at 30 frames per second, the high data volume will create strong needs for data transfer and computation in a real-time framework. Therefore, this work simplifies the process with a lower temporal resolution that still produces accurate speed estimations from a traffic engineering perspective.
In total, SimBarca consists of 101 simulation runs, including one with the original demand profile unmodified and the other 100 with different demands augmented from the original one. Out of all the simulation sessions, 75 runs are randomly selected for training and the rest are for testing. In the graph-structured data, there are 1570 nodes and 2803 edges in total, and the adjacency matrix is a symmetric binary matrix by assuming the graph is undirected. The experiments start with a full-information setting where the model has access to segment speeds and point speeds for all road segments (Section IV-C and IV-E), and then move on to a more realistic scenario where the model only has access to noisy data with partial spatial coverage (Section IV-F).
IV-A Data Visualization
@@
In the first place, the traffic statistics required by Equation 6 and 7, e.g., total distance and time traveled by all vehicles, are calculated at each simulation time step (0.5s), and then utilized for speed computation. Figure 6 compares the segment speed and point speed per simulation time step, which are the finest possible details from the simulation data. The segment speed is computed with all the vehicle trajectories in the road segment, and therefore it can reflect the overall traffic condition in the segment. There is a strong periodic pattern in segment speed due to traffic lights, with which one can identify the acceleration and deceleration patterns in the segment speed. On the other hand, the point speed can only capture the instant speed of vehicles at a specific location, and thus contains more stochastics due to randomized vehicle passing. As a result, from the aspect of congestion development and propagation, point speed does not provide as much information as the segment speed. Besides, the randomized simulations have resulted in varying dynamics in the transportation network. For example, Figure 7 shows the segment speed of the same road segment from multiple simulation runs. The periodic pattern is consistent across different runs, but the details vary because of the demand augmentation and vehicle generation processes in the simulation.
@@
Figure 6: Comparison of segment speed and point speed per simulation time step (0.5s), segment ID 9971
@@
Figure 7: segment speed from multiple simulation runs, segment ID 9971
@@
The initial statistics are then aggregated to coarser time resolutions for the training of the prediction model. In all the experiments, the model is trained to forecast the future segment speed rather than future point speed, as the former can better reflect the traffic dynamics in the whole road segment. The differences between point speed and segment speed are still clear in the aggregated data. For example, Figure 8a shows the segment speed and point speed for the road segment with ID 9971 (as marked in Figure 3), both aggregated every 3 minutes. In this segment, point speed overestimates true segment speed roughly by 5 m/s.
@@
Figure 8b shows an example training sample obtained by the sliding window approach. In the input sequences (first 30 minutes), the segment speed (Drone Input) is given every 5 seconds, where the periodic pattern can be visually identified as the per-time-step case in Figure 6. For example, the segment speed drops to zero periodically at every cycle of traffic signals due to the effect of the red phase. The point speed (LD Input) is aggregated to a 3-minute interval as common practice, and still it overestimates the true speed, since the measured loop detector speed is influenced by the distance of the sensor from the stop line. The second 30 minutes of the figure shows the prediction labels, including the segment-level future speed (Segment Label) as well as the future speed of the region (Regional Label) that this segment belongs to. In this example, traffic congestion has already built up within the short time span of the training sample, as both the segment and regional speed have decreased.
@@
Figure 9: Regions for regional prediction and grids for drone monitoring
Figure 10 shows the Macroscopic Fundamental Diagram (MFD) of the whole road network under different demand scales, which is the augmentation factor multiplied with the demand matrix as mentioned in Section III-C. Notably, the exact number of vehicles in the simulation is still affected by the other two augmentation methods, and thus is not strictly proportional to the demand scale. In these plots, the vehicle flow is computed by $q=D/tL$, where $D$ is the total distance traveled by all vehicles during the time interval of concern $t$, and $L$ is the total length of the road network. Similarly, the vehicle density is obtained using $k=T/tL$ with $T$ being the total travel time of all vehicles. Each subplot shows the MFD with $t$ being 5 seconds and 3 minutes, corresponding to the time resolution of drone and loop detector data respectively, and the data points are annotated with the time from the simulation start.
Figure 10: MFD under different demand scales, (a) 120%, (b) 150% and (c) 180%.
These plots show that as the demand level increases, the vehicle flow becomes saturated earlier and the vehicle density reaches a higher maximum value. Meanwhile, the network-level average speed also decreases greatly as the congestion becomes more severe. Besides, the hysteresis loop is consistent with previous empirical and simulation observations [50]. The increased congestion can be further evidenced by the travel time distribution of the network, as shown in Figure 11. The increase in demand scale leads to a larger average and a longer tail in the travel time distribution, which means more and more vehicles are experiencing longer travel times.
@@
IV-B Implementation Details
To provide a reasonable baseline model, the implementation of HiMSNet is kept as concise as possible. The first step in HiMSNet is value embedding, i.e., the values and corresponding timestamps are embedded into a feature space using a linear layer, whose output dimension is 64. Specifically, the missing values, which appear when a road is not monitored or no vehicle passes a monitored road, are replaced with learnable neural network parameters of equal size. After the embedding, the drone modality has 2 layers of 1D convolution with kernel size 3 and stride 3 to down-sample the temporal resolution. Then, each modality has an individual temporal encoder $\mathcal{H}_{e}$ consisting of 3-layer LSTMs with hidden size 64, and their last step outputs will be used as the local feature $\mathbf{z}_{p}$. The decoders for both segment-level and regional predictions ($\mathcal{H}_{d}$ and $\mathcal{H}_{I}$) are 2-layer MLPs with doubled hidden size (128) and output dimension 10 to predict 10 future time steps (30 minutes). Contrary to modality-specific temporal encoders, the decoder is shared as the features have been combined. The message encoding layer $\mathcal{R}_{1}$ is a 1 $\times$ 1 convolution layer that projects the features while keeping the dimension unchanged. Accordingly, the message decoding layer $\mathcal{R}_{2}$ inherits the same structure. The GME module is implemented with 3 layers of GCNConv [51] followed by ReLU activation and Layer Normalization [52].
In the experiments, Mean Absolute Error (MAE) is used as the loss function for both segment-level and regional prediction tasks, and equal weights are assigned to the two tasks by default. HiMSNet is trained for 30 epochs using Adam optimizer with a learning rate 0.001, weight decay $1\times 10^{-4}$, $\beta=(0.9,0.999)$ and batch size 8. We use a learning rate scheduler that gradually increases the learning rate from 0 to the chosen value during the first epoch, and decreases it to 10% and 1% of the chosen value at 70% and 85% of the total epochs. All implementations are based on PyTorch, and the experiments are conducted on a computer equipped with an Nvidia RTX 4090 GPU.
IV-C Overall Prediction Performance
@@
Table I compares the prediction errors of HiMSNet with three simple baselines, best results in bold. Last Observation (LO) uses the last observed speed as the segment-level prediction, while Input Average (IA) uses the average speed of the input sequence. Both of them can be configured to use loop detector (ld) or drone (drone) data, and the regional prediction is made by averaging the segment-level predictions. The Label Average (LA) baseline uses the average speed of the whole test set labels as prediction, which applies to both prediction tasks. In fact, the LA baseline is the best possible constant-value prediction that leads to the lowest MAE. The second group of methods are variants of HiMSNet, where we present a brief ablation study to examine the importance of different components. The exact settings are: disable the GME module (/gnn) to exclude vicinity information, only use loop detector modality in the input (ld), and only use the drone modality (drone) in the input.
The IA and LO baseline with loop detector data have significantly higher errors than the variants using drone data, which originates from the difference between loop detector measurements and the drone measurements as shown in Figure 6. IA(drone) also outperforms LO(drone), which is expected as the drone-measured speed has a high update frequency (every 5s) to reveal the traffic dynamics within a signal cycle. Therefore using the last observation as a prediction will be very biased towards the end of the cycle. The LA baseline has the lowest error among the first group of results, and HiMSNet outperforms all the baselines, especially in the regional prediction task.
TABLE I: Prediction errors for road segment task and regional task
@@
For the regional task, most of the HiMSNet variants have close prediction performance, except that HiMSNet(ld) clearly has higher errors than the other variants. Moreover, although the drone modality has been excluded from the input of HiMSNet(ld), the training labels of this model are still computed using the vehicle trajectory data, and in reality, such data are not possible with loop detectors only. Compared to using drone data directly as input, HiMSNet(ld) faces additional challenges because of the difference between point speeds and segment speed (Figure 6). This performance difference suggests that the translation from point speed to segment speed can not be easily learned, and also highlights the necessity of drone data modality.
@@ IV-D Detailed Analysis of Prediction Performance
Figure 14: Average MAE (unit m/s) over all prediction horizons grouped by different factors. (a) by demand scale for the segment-level task, (b) by demand scale for the regional task, (c) by average segment speed over the test set (segment-level task).
@@ IV-E Effects of Hyperparameters
The choices of hyperparameters can affect the performance of the model. Figure 15 compares the 30-min-ahead prediction MAE under different hyperparameter settings, including adjacency hops, loss balance between tasks, training epochs and model sizes. The ablation study in Table II has shown that vicinity information is important for the segment-level prediction task, which highlights the effectiveness of the GME module. Still, the scope of the considered neighborhood should be properly chosen. In Figure 15a, a single GCNConv layer with 1-hop adjacency will strictly follow the road network definition and only take into account the immediate neighbors of a road segment. Although this message passing scheme is stacked 3 times in the GME module, it still turns out to be suboptimal due to the limited adjacency scope. On the contrary, a layer with 5-hop adjacency will have a large scope and include the nodes that are 5 hops away from a road segment, and thus it may introduce much irrelevant information. Therefore in the three experimented setups, the 3-hop adjacency layer achieves the best performance for the segment-level task, as it has the most suitable vicinity scope.
@@
IV-F Partial Observation and Noisy Data

### v2

@@ III-C Dataset Creation
Although existing drone datasets have already provided high-quality vehicle trajectory data, e.g., pNEUMA [10] and highD [53], the aim of this work is to predict the traffic states for all road segments within a large urban region, which is beyond the scope of available datasets. For this reason, detailed vehicle trajectory data are generated using the microscopic traffic simulator Aimsun and its road network for central Barcelona [54]. Concretely, the dataset creation process consists of three steps: 1) Simulate Vehicle Trajectories, 2) Compute Traffic Variables and 3) Sample Extraction and Sensor Simulation.
III-C1 Simulate Vehicle Trajectories
The simulation environment has been properly calibrated based on real-world traffic scenarios, and it has been applied in various analyses on urban traffic, such as network wide-perimeter control [55, 56]. As shown in Figure 3, the network consists of many road segments, intersections as well as centroids ($\odot$ like) that serve as origins and destinations of vehicles. The fine details from microscopic traffic simulations allow great flexibility for subsequent processing and analysis, e.g., the computation of vehicle flow, density and speed. Concretely, there are mainly three types of available information:
@@ III-C1 Simulate Vehicle Trajectories
III-C2 Compute Traffic Variables
Using vehicle trajectory data from microscopic traffic simulation, this work computes point speed and segment speed to resemble the characteristics of inductive loop detectors and drones. Inductive loop detectors can only measure vehicle speeds at certain points, due to fixed installations. Since vehicles can change their speeds within a road segment, the such speed measurements are valid only in its proximity and often cannot represent the situation of the whole segment. On the contrary, drones can capture nearly all vehicles in their field of view, therefore it is possible to compute accurate segment speeds using all available vehicle trajectories. In this work, all road segments are assumed to have a loop detector at their middle points.
@@ III-C2 Compute Traffic Variables
For a loop detector installed $d$ meters from the start, if a vehicle crosses the detector’s position during the time interval $(t_{s},t_{e})$, its corresponding position will satisfy $x_{s}<d<x_{e}$. Then, the vehicle is detected at the starting timestep of the trajectory split. For example, the vehicle in Figure 4 is detected at $t_{2}$, because its trajectory crosses the loop detector between $t_{2}$ and $t_{3}$. We count the number of detected vehicles $N$ in a time interval (e.g., every 3 minutes), and arithmetically average their speeds to get point speed as
@@
By choosing different time resolutions for Equation 9 and 10, the simulated loop detector and drone measurements can be computed individually for each road segment. Similarly, regional speed can be obtained by applying Equation 9 to all the trajectory splits in a region during a time interval.
III-C3 Sample Extraction and Sensor Simulation
With the segment speed and point speed for each road segment, and average speed for each region, the samples for the prediction model can be generated using a sliding window as in Figure 5.
Each simulation session lasts for 4 hours in the simulation world, with a 15-minute warm-up and 1 hour 45 min main demand, while the real-world run time differs according to the traffic demand. This demand profile is intended to simulate the morning peak hours in the road network. The sliding window will start at the end of the warm-up period up to 2 hours after the beginning of the main demand, and the window size is 1 hour with a 3-minute step size between consecutive samples. Inside a window, the first half is used as input and the second half is used as targeted output for the prediction model. Therefore each simulation session will have 20 samples, each with 30-minute input and 30-minute output.
@@ III-C3 Sample Extraction and Sensor Simulation
TABLE I: Input and target sequences in SimBarca
Speed Sensor Input In-Step Pred. Pred-Step
Segment Drone 5 s 3 min
Point LD 3 min $\times$ -
Regional - $\times$ - 3 min
As shown in Table I, drones and loop detectors are assumed to provide segment speed and point speed respectively. Point speed in the prediction window is excluded from targeted outputs because it can not represent the situation of the whole segment. Meanwhile, regional speed is excluded from input sequences since it is highly-aggregated and can not be directly measured by an individual sensor. Therefore, SimBarca contains two input modalities, i.e., point speed and segment speed from simulated loop detectors and drone, and two prediction tasks, i.e., segment-level and regional speed prediction.
In the input, loop detectors speeds are given every 3 minutes according to literature [58] and drones report speeds every 5 seconds. Although in reality, drones are able to provide high-resolution video at 30 frames per second, the high data volume will create high demand for data transfer and computation in a real-time framework. Therefore, this work simplifies the process with a lower temporal resolution while preserving accurate speed estimations from a traffic engineering perspective. The segment speeds computed by Equation 9 are used as segment-level labels, and the regional labels are obtained by applying the same formula to all trajectory splits within a region. Both labels have 3 minute time step.
@@ IV Experiments
In this part, Section IV-A visualizes the dataset at both micro and macro levels, and Section IV-B introduces the experiment settings for deep learning. Section IV-C, IV-D and IV-E start the evaluation with a full-information setting where the model has access to segment speeds and point speeds for all road segments. Then, Section IV-F and IV-G study a more realistic scenario with noisy data and partial sensor coverage.
IV-A Visualization of Fine-grained and Aggregated Data
@@
Figure 6 compares the segment speed and point speed per simulation time step, which are the finest possible details from the simulation data. Traffic lights have caused a strong periodic pattern in segment speed, i.e., the acceleration and deceleration patterns in the segment speed. On the other hand, the point speed can only capture the instant speed of vehicles at a specific location, and thus contains more stochastics due to randomized vehicle passing. From the aspect of congestion development and propagation, point speed does not provide as much information as the segment speed. Besides, the randomized simulations have resulted in varying dynamics in the transportation network. For example, Figure 7 shows the segment speed of the same road segment from multiple simulation runs. The periodic pattern is consistent across different runs, but the details vary because of the randomized demand augmentation and vehicle generation processes.
@@
Figure 6: Comparison of segment speed and point speed per simulation time step (0.5s), segment ID 9971. This visualization does not include missing values when a loop detector or a drone senses no vehicle.
@@
Figure 7: segment speed from multiple simulation runs per simulation time step (0.5s), segment ID 9971. The time series are drawn as discrete points to reduce overlapping.
@@
The fine-grained initial statistics are then aggregated to coarser time resolutions for the deep learning experiments. The differences between point speed and segment speed are still clear in the aggregated data. For example, Figure 8a shows the segment speed and point speed for the road segment with ID 9971 (as marked in Figure 3), both aggregated every 3 minutes. In this example, point speed overestimates true segment speed roughly by 5 m/s, because a loop detector can not sense the vehicles stopping to wait for the red lights unless the queue is long enough to reach it.
@@
Figure 8b shows an example training sample obtained by the sliding window approach. In the input sequences (first 30 minutes), the segment speed (Drone Input) is given every 5 seconds, where the periodic pattern can be visually identified as the per-time-step case in Figure 6. For example, the segment speed drops to zero periodically at every cycle of traffic signals due to the effect of the red phase. The point speed (LD Input) is aggregated to a 3-minute interval as common practice, and still it overestimates the true speed, since the measured loop detector speed is influenced by the distance of the sensor from the stop line. The second 30 minutes of the figure shows the prediction labels, including the future speed of the segment (Segment Label) and that of the region (Regional Label) this segment belongs to. In this example, traffic congestion has already built up within the short time span of the training sample, as both the segment and regional speed have decreased.
@@
Figure 9: Spatial regions and grids for drone monitoring.
Figure 10 shows the MFD of the whole road network under different demand scales, which is the augmentation factor multiplied with the demand matrix as mentioned in Section III-C1. Notably, the exact number of vehicles in the simulation is still affected by the other two augmentations, and thus is not strictly proportional to the demand scale. The vehicle flow is obtained by $q=D/tL$, where $D$ is the total distance traveled by all vehicles during the time interval of concern $t$, and $L$ is the total length of the road network. Similarly, the vehicle density is obtained using $k=T/tL$ with $T$ being the total travel time of all vehicles. Each subplot shows the MFD with $t$ being 5 seconds and 3 minutes, corresponding to the time resolution of drone and loop detector data respectively, and the data points are annotated with the time from the simulation start.
Figure 10: MFD under different demand scales, (a) 120%, (b) 150% and (c) 180%. The points in the plots are pairs of vehicle flow and density values, linked according to simulation time indicated by the text annotations. These MFDs have hysteresis because after the duration with traffic demand, the simulation still runs to allow vehicles to clear the network. This exiting stage behaves differently due to the absence of new vehicles, and it is not included in sample extraction, as shown in Figure 5.
These plots show that as the demand level increases, the vehicle flow becomes more saturated and the vehicle density reaches a higher maximum value. Meanwhile, the network-level average speed also decreases greatly as the congestion becomes more severe. Besides, the hysteresis loop is consistent with previous empirical and simulation observations [60]. The increased congestion can be further evidenced by the travel time distribution of the network, as shown in Figure 11. The increase in demand scale leads to a larger average and a longer tail in the travel time distribution, which means more and more vehicles are experiencing longer travel times.
@@
IV-B Experiment and Evaluation Settings
SimBarca consists of 101 simulation runs, including one with the original demand profile unmodified, provided along side the simulation network by Aimsun [54], and the other 100 containing different demands augmented from the original one. The training set contains 75 randomly selected simulation runs, and the rest are for testing. The data graph is undirected and contains 1570 nodes (road segments) and 2803 edges.
By default, HiMSNet is trained for 30 epochs using Adam optimizer with a learning rate 0.001, weight decay $1\times 10^{-4}$, $\beta=(0.9,0.999)$ and batch size 8. The learning rate scheduler gradually increases the learning rate from 0 to the chosen value during the first epoch, and decreases it to 10% and 1% of the chosen value at 70% and 85% of the total epochs. All implementations are based on PyTorch, and the experiments are done on a computer with an Nvidia RTX 4090 GPU.
@@ IV-B Experiment and Evaluation Settings
IV-C Overall Prediction Performance
Table II compares the prediction errors of HiMSNet with three simple baselines to demonstrate the necessity of drone data and the effect of HiMSNet modules, best results in bold.
Last Observation (LO) uses the last observed speed as the segment-level prediction, while Input Average (IA) uses the average speed of the input sequence. Both of them can be configured to use loop detector (ld) or drone (drone) data, and the regional prediction is made by averaging the segment-level predictions. The Label Average (LA) baseline uses the average speed of the whole test set labels as prediction, which applies to both prediction tasks. In fact, the LA baseline is the best possible constant-value prediction that leads to the lowest MAE. The second group of methods are variants of HiMSNet, where we present a brief ablation study to examine the importance of different components. The exact settings are: replacing the lightweight GCNConv [51] with a Transformer Encoder (TF) [33], disable the GME module (/gnn) to exclude vicinity information, only use loop detector modality in the input (ld), and only use the drone modality (drone) in the input.
The IA and LO baseline with loop detector data have significantly higher errors than the variants using drone data, which originates from the difference between loop detector measurements and the drone measurements as shown in Figure 6. IA(drone) also outperforms LO(drone), which is expected as the drone-measured speed has a high update frequency (every 5s) to reveal the traffic dynamics within a signal cycle. Therefore using the last observation as a prediction will be very biased towards the end of the cycle. The LA baseline has the lowest error among the first group of results, and HiMSNet outperforms all the baselines, especially in the regional prediction task. Transformer [33] is generally considered as a more advanced architecture, however, since the HiMSNet(TF), has very similar errors as the HiMSNet with GCNConv [51], we default to the simpler one.
TABLE II: Prediction errors for road segment task and regional task in perfect information case
@@ IV-C Overall Prediction Performance
For the regional task, most of the HiMSNet variants have close prediction performance, except that HiMSNet(ld) clearly has higher errors than the other variants. Moreover, although the drone modality has been excluded from the input of HiMSNet(ld), the training labels of this model are still computed using the vehicle trajectory data, and in reality, such data are not possible with loop detectors only. Compared to using drone data directly as input, HiMSNet(ld) faces additional challenges because of the difference between point speeds and segment speed (Figure 6). This performance difference suggests that the translation from point speed to segment speed can not be easily learned, and also highlights the necessity of drone data.
@@ IV-D Detailed Analysis of Prediction Performance
This section visualize the model’s predictions, as well as the distribution of prediction errors by spatial locations, demand scales and average speeds.
@@
Figure 14: Average MAE (unit m/s) over all prediction horizons grouped by different factors. (a) by demand scale for the segment-level task, (b) by demand scale for the regional task, (c) by average segment speed in the test set (segment-level task).
@@ IV-E Effects of Hyperparameters
This section compares the performance of the model under different hyper-parameters to justify the choice of default training settings. Figure 15 compares the 30-min-ahead prediction MAE under different hyperparameter settings, including adjacency hops, loss balance between tasks, training epochs and model sizes. The ablation study in Table III has shown that vicinity information is important for the segment-level prediction task, which highlights the effectiveness of the GME module. Still, the scope of the considered neighborhood should be properly chosen. In Figure 15a, a single GCNConv [51] layer with 1-hop adjacency will strictly follow the road network definition and only take into account the immediate neighbors of a road segment. Although this message passing scheme is stacked 3 times in the GME module, it still turns out to be suboptimal due to the limited adjacency scope. On the contrary, a layer with 5-hop adjacency will have a large scope and include the nodes that are 5 hops away from a road segment, and thus it may introduce much irrelevant information. Therefore in the three experimented setups, the 3-hop adjacency layer achieves the best performance for the segment-level task, as it has the most suitable vicinity scope.
@@
IV-F Experiment Settings with Partial and Noisy Data